# Digital Simulation of Non-Abelian Parafermions in Superconducting Circuits

### v4

@@ Abstract
Parafermions, which can be viewed as a fractionalized version of Majorana modes, exhibit profound non-Abelian statistics and emerge in topologically ordered systems, while their realization in experiment has been challenging. Here we propose an experimental scheme for the digital simulation of parafermions and their non-Abelian braiding statistics in superconducting (SC) circuits by realizing the $\mathbb{Z}_{d}$ plaquette model on a two-dimensional lattice. Two protocols using quantum circuits and non-destructive measurements are introduced to prepare the ground state, on which parafermion pairs are created by engineering dislocations. We then propose a generalized code deformation approach to realize the fusion and non-Abelian braiding of parafermion modes, and show the application of this approach to the $\mathbb{Z}_{3}$ parafermions. We also examine the experimental parameter regime to confirm the feasibility of our scheme in SC devices. This work extends previous quantum simulation of twist defects in SC qubits to qudit systems, and opens up a way for parafermion-based high-dimensional topological quantum computing with experimental feasibility.
@@ I Introduction
The recent rapid advancement in quantum processors enables simulation of many-body quantum states and exploration of the intriguing physics they exhibit [1, 2, 3, 4, 5, 6]. This provides a way to study and manipulate exotic particles that have yet to be realized or are difficult to manipulate in real materials [7, 8, 9, 10, 11]. A notable example is the simulation of anyons and topological defects in topologically ordered states (topological orders) [3, 4, 12, 13, 14, 15, 16, 17]. In particular, recent progress in digital simulation of twist defects in the $\mathbb{Z}_{2}$ toric code model with superconducting (SC) qubits demonstrates the (projective) non-Abelian braiding statistics [3, 4]. These efforts advance topologically protected quantum information processing, providing a foundation for the potential realization of topological quantum computing.
@@
Twist defects in the $\mathbb{Z}_{2}$ toric code model exhibit non-Abelian statistics of Majorana fermion modes, which is described by the Ising anyon model [18, 19]. Quantum computing schemes based on Majorana modes have been widely studied, but their experimental realization in one- or two-dimensional condensed matter systems faces significant challenges [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48]. More critically, braiding of Majorana modes alone is insufficient for universal quantum computing [49, 50]. Universal quantum computing schemes based on Majorana modes require additional non-topological assistance methods, making them not strictly topological [51, 52]. $\mathbb{Z}_{d}$ parafermions generalize Majorana modes (which are a special case for $d=2$) and exhibit more complex and nonlinear non-Abelian braiding statistics [53, 54]. In contrast to Majorana modes, parafermions can encode multi-level computational units, or qudits, which offer a larger state space for storing and processing information, as well as advantages in reducing circuit complexity and improving algorithm efficiency [55]. Braiding of parafermions allows the generation of the single-qudit Clifford group for any $d$, and the many-qudit Clifford group for odd $d$ [56]. Fault-tolerant non-Clifford gates can be performed by magic state distillation with high fidelity [57, 58, 59].
@@
Proposals for realizing parafermions typically involve systems with strong electron-electron interactions, which however pose serious experimental challenges [60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75]. On the other hand, to engineer the topological defects is an alternative general approach for realizing non-Abelian statistics within Abelian topological orders [76]. This provides a simpler way to realize parafermion defects in Abelian topologically ordered states. Furthermore, the braiding of Abelian anyons of the underlying state facilitates performing topological non-Clifford gates. A typical class of stabilizer Hamiltonian models that realize Abelian topological orders is the $\mathbb{Z}_{d}$ plaquette models with qudits on a two-dimensional square lattice, in which certain lattice dislocations realize parafermions [77]. As a stabilizer model, the digital simulation of its ground state and excited states with a few quasiparticles is in principle achievable on experimental platforms with coherent control of qudits. The SC devices, such as transmons, which possess readily addressable higher energy levels, constitute promising candidates for implementing such platforms [78]. Recent works on SC devices have demonstrated high-fidelity single- and two-qutrit (qudit with $d=3$) entangling operations [79, 80, 81, 82, 83, 84, 85]. Nevertheless, realization of the parafermion modes and implementation of their non-Abelian statistics necessitate qualitatively more complicated quantum gate operators and involve addressing non-trivial physical challenges in comparison with the Majorana modes [4, 3], and it remains unclear how to achieve such quantum simulation in the real SC circuits. More precisely, while the generalization from Majorana mdoes to parafermion modes may appear straightforward at the level of mathematical model construction, its physical realization remains significantly non-trivial. First, the fundamental single- and two-qudit gates are unitary but not Hermitian, requiring careful consideration of both the operations and their Hermitian conjugates in circuit implementations. Second, the multi-level structure of qudits necessitates novel physical processes to control level transitions and to implement entangling operations involving higher levels. Examples include the two-photon process for controlling the 00-$2$ transition in qutrits [82] to realize high-fidelity generalized Hadamard gate, and implementing control gates based on tunable cross-Kerr interactions [81, 84]. Third, parafermions exhibit richer non-Abelian braiding statistics and fusion rules compared to Majorana modes, making it physically non-trivial to identify suitable observables for characterizing their braiding and fusion properties. This work aims to address these challenges and integrate recently developed qudit techniques in SC devices to develop an experimentally feasible scheme for simulating parafermions and their fusion and braiding, providing insights into the parafermion-based high-dimensional topological quantum computing.
@@
Here, we summarize the main results of our work. In this work, we propose an experimentally feasible scheme to simulate parafermions and their fusion and non-Abelian braiding in SC circuits, focusing on the $\mathbb{Z}_{d}$ plaquette model on a planar surface. We introduce two efficient ground state preparation protocols: one using quantum circuits and the other employing non-destructive measurements. By engineering dislocations in the lattice, pairs of parafermions are created. With a novel generalized code deformation approach, as proposed in the present work, these parafermions are moved on the planar surface, enabling a feasible way for their fusion and braiding operations. The concrete example of $d=3$ parafermions is studied in detail. The real parameter conditions are examined, with the feasibility of our scheme in experiment being confirmed. We note that while the present simulation is designed for the SC devices, it is also compatible with other experimental platforms capable of implementing the gates that we use. Our work provides a theoretical foundation for the imminent digital simulation of parafermions in quantum computing platforms, and shall advance topologically protected high-dimensional quantum information processing.
@@
The remainder of the paper is organized as follows. In Sec. II, we introduce the $\mathbb{Z}_{d}$ plaquette model, its elementary excitations, and parafermion defects. In Sec. III, we present the protocol for ground state preparation and parafermion creation. Sections IV and V outline the schemes for realizing the fusion and non-Abelian braiding of parafermions. In Sec. VI, we discuss the experimental parameters of the SC devices for realizing our scheme. Finally, in Sec. VII, we summarize our findings and discuss potential applications in high-dimensional topological quantum computing.
@@ II The Model
Figure 1: Illustration of the model with a pair of prafermions. (a) The $\mathbb{Z}_{d}$ plaquette model on a checkerboard lattice with a dislocation colored by light blue, where the small white dots represent qudits. The red cross marks the defect associated with the dislocation, identified as the parafermion mode $\gamma$. The red dotted line indicates the branch cut connecting the two parafermions. The arrowed dark blue curves represent string operators that create $e-\overline{e}$ and $m-\overline{m}$ excitations, which reside on the light and dark plaquettes, respectively. In our convention, the $e$ or $m$ particle is located at the start of the string operator and its dual is located at the end of the string operator. Stabilizers $O_{p}$ on both dark and light plaquettes are shown. Graphical representations of the generalized Pauli operators, commutation relation, and stabilizers are shown. (b) An illustration of the energy diagram and microwave driving of a typical transmon as a multilevel system. The frequency difference $\omega_{12}-\omega_{01}$ characterizes the anharmonicity. (c) Dislocations are engineered with two neighboring stabilizers being replaced with a single one. Applying a two-qudit quantum circuit moves and separates the parafermion defects. (d) Schematic of the cross-Kerr interaction that is used to realize two-qutrit gates. (e) A configuration that realizes the tunable cross-Kerr interaction using a frequency-tunable coupler, as demonstrated in Ref. [81].
@@
where the labels of the qudits are shown in the top of Fig. 1(a), and the sum of plaquettes $p$ is for all the light and dark plaquettes. Here, $Z$, $X$, $Z^{\dagger}$, and $X^{\dagger}$ are generalized Pauli operators for qudits. They are defined by the properties: $X^{d}=\mathbbm{1},\,Z^{d}=\mathbbm{1},\,XZ=\omega ZX,$ where $\omega=e^{\frac{2\pi i}{d}}$ and act on the qudit basis as $Z\left|{m}\right\rangle=\omega^{m}\left|{m}\right\rangle$ and $X\left|{m}\right\rangle=\left|{m-1\,(\text{mod}\quad d)}\right\rangle$. The experimental realization of single-qudit operations can be achieved through microwave driving resonant with specific qudit state subspaces. For $d=3$ (i.e., a qutrit), any single-qutrit gate can be decomposed into a sequence of operations acting on two-level subspaces [86, 83], and implemented with high fidelity by accounting for the AC Stark effect arising from anharmonicity [86, 87]. It is convenient to introduce a graphical representation [Fig. 1(a)] for the generalized Pauli operators and their commutation relations. With this representation, the stabilizer $O_{p}$ is described by a loop operator around the plaquette $p$, which satisfies $O_{p}^{d}=1$ and has eigenvalues $1$, $\omega$,…, $\omega^{d-1}$. Stabilizers on different plaquettes commute with each other; therefore, the Hamiltonian Eq. (1) is exactly solvable. The ground state $\left|{\Psi_{0}}\right\rangle$ satisfies:
@@
The model defined by Eq. (1) provides a Hamiltonian realization of the topological gauge theory with gauge group $\mathbb{Z}_{d}$. The stabilizers on light (dark) plaquettes correspond to local Gauss’s law (flatness condition) constraints. The ground state, defined by Eq. (2), is the state where all local constraints are satisfied, while excited states arise from violations of these constraints. For $d=2$, the stabilizer has only two eigenvalues $1$ and $-1$, indicating whether a local constraint is satisfied or violated. The elementary excitations—violations of Gauss’s law and flatness condition—are termed charge ($e$) and flux ($m$) excitations, respectively, following its physical meaning in gauge theory. In the $\mathbb{Z}_{2}$ gauge theory, these excitations are self-dual, satisfying $e^{2}=1$ and $m^{2}=1$. For $d\geq 3$, however, charge and flux excitations are no longer self-dual and can appear in multiple powers. Specifically, the elementary excitations of Eq. (1) include charge excitations $e,e^{2},...,e^{d-1}$ on light plaquettes and flux excitations $m,m^{2},...,m^{d-1}$ on dark plaquettes [see Appendix. A for an algebraic characterization of these excitations]. More general excitations involve composites $e^{k}m^{d-k}$, known as dyons. When a charge excitation $e$ encircles a flux excitation $m$ counterclockwise, or vice versa, the system acquires a phase factor $\overline{\omega}\equiv\omega^{d-1}$ due to the Aharonov-Bohm effect. Consequently, charge, flux, and dyon excitations are Abelian anyons, created by string operators [graphically represented in Fig. 1(a)]. The ground state of Eq. (1) exhibits the Abelian topological order characterized by the quantum double $D(\mathbb{Z}_{d})$ [88].

### v5

# Simulating non-Abelian statistics of parafermions with superconducting processor

@@ Abstract
Parafermions, which can be viewed as a fractionalized version of Majorana modes, exhibit profound non-Abelian statistics and emerge in topologically ordered systems, while their realization in experiment has been challenging. Here we propose a novel experimental scheme for the quantum simulation of parafermions and their non-Abelian braiding statistics in superconducting (SC) circuits by realizing the $\mathbb{Z}_{d}$ plaquette model on a two-dimensional lattice. Two protocols using quantum circuits and non-destructive measurements are introduced to prepare the topologically ordered ground state, on which parafermion pairs are created by engineering dislocations. We then propose a generalized code deformation approach to realize the fusion and non-Abelian braiding of parafermions, and show the application of this approach to the $\mathbb{Z}_{3}$ parafermions. We also examine the real experimental parameter regime to confirm the feasibility of our scheme in SC devices. This work extends previous quantum simulations of topological defects in SC qubits to qudit systems and opens up a promising way for parafermion-based high-dimensional topological quantum computing with experimental feasibility.
@@ I Introduction
The rapid advancement of quantum processors has enabled the simulation of many-body quantum states and the exploration of intriguing quantum phenomena [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. An particularly exciting direction is the simulation of anyons [25, 26, 27, 28, 29, 30], exotic quasiparticles emerging in topologically ordered states (topological orders) [31, 32], whose realization and manipulation remain challenging in materials [33, 34, 35, 36, 37, 38]. Recent digital simulations using superconducting (SC) qubits have demonstrated non-Abelian braiding statistics of engineered topological defects in the $\mathbb{Z}_{2}$ toric code model [39, 40, 41, 42, 43, 44]. These defects, associated with one-dimensional lattice dislocations, exhibit fusion and projective braiding properties identical to those of intrinsic point-like non-Abelian anyonic excitations [45], enabling the global encoding of quantum information [46]. Consequently, braiding these topological defects to process quantum information is topologically protected at the logical level [47]. This protection, rooted in the topological nature of the surface code model and the non-local encoding with topological defects, offers a promising pathway toward fault-tolerant quantum computing [48, 49, 42, 43, 50, 51, 52, 53].
@@
Topological defects in the $\mathbb{Z}_{2}$ toric code model manifest the non-Abelian statistics of Majorana fermion modes, which is described by the Ising anyon model [40, 54]. Quantum computing schemes based on Majorana modes have been widely studied [55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82], but their experimental realization in one- or two-dimensional condensed matter systems faces significant challenges [83, 84, 85, 86, 87, 88, 89, 90, 91, 92]. More critically, braiding Majorana modes alone is insufficient for universal quantum computing [93, 94]. Achieving universality requires additional non-topological assistance methods, which compromise the inherent topological protection [95, 96, 97]. $\mathbb{Z}_{d}$ parafermions generalize Majorana modes (which are a special case for $d=2$) and exhibit more complex and nonlinear non-Abelian braiding statistics [98, 99]. In contrast to Majorana modes, parafermions can encode multi-level computational units, known as qudits, which provide a larger state space for storing and processing information [100, 101, 102]. This enables the reduction in circuit complexity and enhancement in algorithmic efficiency [103]. High-dimensional quantum computing with qudits also offers further advantages, such as greater flexibility in simulating quantum dynamics [104, 100] and higher-fidelity magic state distillation [105, 106]. Unlike Majorana modes, braiding of parafermions enables entanglement gates. In particular, it allows the generation of the single-qudit Clifford group for any $d$ and the multi-qudit Clifford group for odd $d$ [107]. Non-Clifford gates can be implemented by incorporating flux braiding through the Aharonov–Casher effect [108], complementing parafermion braiding to realize a universal gate set. Moreover, parafermions can serve as building blocks for engineering Fibonacci anyons [109, 110], whose braiding supports universal quantum computation. As such, $\mathbb{Z}_{d}$ parafermions with $d>2$ are computationally more powerful than Majorana modes, bringing us closer to universal quantum computing.
Proposals for realizing parafermions in condensed matter systems—such as fractional quantum Hall states or fractional topological insulators with induced superconductivity, and coupled nanowires—typically require strong interactions or fractionalization, which however pose significant experimental challenges [111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128]. Alternatively, engineering topological defects in Abelian topological orders provides a general route to realizing non-Abelian statistics [45], which can extend beyond the Majorana case. Since Abelian topological orders are more prevalent and generally easier to realize than their non-Abelian counterparts, this approach offers a simpler and more experimentally feasible path to realizing parafermions. Additionally, existing Abelian anyonic excitations facilitate the implementation of phase gates. Following this idea, topological defects obeying parafermion statistics are theoretically predicted in a stabilizer Hamiltonian model known as the $\mathbb{Z}_{d}$ plaquette model [129], though experimental realizations remain limited to the simplest case of Majorana modes ($d=2$[130, 42, 43]. Digital quantum simulation of the $\mathbb{Z}_{d}$ plaquette model with topological defects presents a promising approach for realizing parafermions, enabling state preparation and evolution via quantum circuits without relying on microscopic Hamiltonians [129, 131, 132, 133, 51]. Well-separated topological defects can globally encode logical qudits, functioning similarly to point-like intrinsic parafermionic excitations [40, 134]. In practical implementations, an error correction strategy is needed to manage stray anyons [135]. Fortunately, parafermions realized through defects associated with extended structures exhibit enhanced robustness due to their extrinsic nature (i.e., they are not part of the spectrum of the underlying Abelian model) and limited mobility [136, 137]. As a result, non-Abelian error correction is not required; standard surface code error correction on the background Abelian model is sufficient to ensure fault tolerance [138]. In surface code error correction, topological protection is characterized by the code distance—the number of local operations required to transform one ground state into another. For the $\mathbb{Z}_{d}$ plaquette model with parafermion defects positioned at the center of a planar surface, the code distance grows linearly with system size, resulting in an exponential suppression of logical qudit errors. This approach is particularly relevant for noisy intermediate-scale quantum (NISQ) devices [139], offering valuable insights into fault-tolerant, high-dimensional quantum information processing.
@@
Simulating the $\mathbb{Z}_{d}$ plaquette model with parafermion defects requires precise coherent control of qudits. SC devices, such as transmons with readily addressable higher energy levels, offer a promising platform for this purpose [140]. Recent advancements have demonstrated high-fidelity single- and two-qutrit (qudit with $d=3$) entangling operations  [141, 142, 143, 144, 145, 146, 147]. Nevertheless, realization of the parafermion modes and implementation of their non-Abelian statistics necessitate qualitatively more complicated quantum gate operators and involve addressing non-trivial physical challenges in comparison with the Majorana modes [43, 42], and it remains unclear how to achieve such quantum simulation in the real SC circuits. While the mathematical extension from Majorana to parafermion modes may seem straightforward, physical realization is considerably more intricate. First, the fundamental single- and two-qudit gates are unitary but not Hermitian, requiring careful consideration of both the operations and their Hermitian conjugates in circuit implementations. Second, the multi-level structure of qudits necessitates novel physical processes to control level transitions and to implement entangling operations involving higher levels. Examples include the two-photon process for controlling the 00-$2$ transition in qutrits [144] to realize high-fidelity generalized Hadamard gate, and implementing control gates based on tunable cross-Kerr interactions [143, 146]. Third, parafermions exhibit much richer non-Abelian braiding statistics and fusion rules compared to Majorana modes, making it physically non-trivial to identify suitable observables for characterizing their braiding and fusion properties. This work aims to address these challenges and integrate recently developed qudit techniques in SC devices to develop an experimentally feasible scheme for simulating parafermions and their fusion and braiding.
@@
Here, we summarize the main results of our work. In this work, we propose an experimentally feasible scheme to simulate parafermions and their fusion and non-Abelian braiding in SC circuits, focusing on the $\mathbb{Z}_{d}$ plaquette model on a planar surface. We introduce two efficient ground state preparation protocols: one using quantum circuits and the other employing non-destructive measurements. By engineering dislocations in the lattice, pairs of parafermions are created. With a novel generalized code deformation approach, as proposed in the present work, these parafermions are moved on the planar surface, enabling a feasible way for their fusion and braiding operations. The concrete example of $d=3$ parafermions is studied in detail. The real parameter conditions are examined, with the feasibility of our scheme in experiment being confirmed. We note that while the present simulation is designed for the SC devices, it is also compatible with other experimental platforms capable of implementing the gates that we use. Our work provides a theoretical foundation for the imminent quantum simulation of parafermions in quantum computing platforms and advances topologically protected high-dimensional quantum computing.
@@
The remainder of the paper is organized as follows. In Sec. II, we review the $\mathbb{Z}_{d}$ plaquette model, its elementary excitations, and parafermion defects. In Sec. III, we present the protocol for ground state preparation and parafermion creation. Sections IV and V outline the schemes for realizing the fusion and non-Abelian braiding of parafermions. In Sec. VI, we discuss the experimental parameters of the SC devices for realizing our scheme. Finally, in Sec. VII, we summarize our findings and discuss potential applications in high-dimensional topological quantum computing.
@@ II The Model
Figure 1: Illustration of the model with a pair of prafermions. (a) The $\mathbb{Z}_{d}$ plaquette model on a checkerboard lattice with a dislocation colored by light blue, where the small white dots represent qudits. The red cross marks the defect associated with the dislocation, identified as the parafermion mode $\gamma$. The red dotted line indicates the branch cut connecting the two parafermions. The arrowed dark blue curves represent string operators that create $e-\overline{e}$ and $m-\overline{m}$ excitations, which reside on the light and dark plaquettes, respectively. In our convention, the $e$ or $m$ particle is located at the start of the string operator and its dual is located at the end of the string operator. Stabilizers $O_{p}$ on both dark and light plaquettes are shown. Graphical representations of the generalized Pauli operators, commutation relation, and stabilizers are shown. (b) An illustration of the energy diagram and microwave driving of a typical transmon as a multilevel system. The frequency difference $\omega_{12}-\omega_{01}$ characterizes the anharmonicity. (c) Dislocations are engineered with two neighboring stabilizers being replaced with a single one. Applying a certain two-qudit quantum circuit moves and separates the parafermion defects. (d) Schematic of the cross-Kerr interaction that is used to realize two-qutrit gates. (e) A configuration that realizes the tunable cross-Kerr interaction using a frequency-tunable coupler, as demonstrated in Ref. [143].
@@
where the labels of the qudits are shown in the top of Fig. 1(a), and the sum of plaquettes $p$ is for all the light and dark plaquettes. Here, $Z$, $X$, $Z^{\dagger}$, and $X^{\dagger}$ are generalized Pauli operators for qudits. They are defined by the properties: $X^{d}=\mathbbm{1},\,Z^{d}=\mathbbm{1},\,XZ=\omega ZX,$ where $\omega=e^{\frac{2\pi i}{d}}$ and act on the qudit basis as $Z\left|{m}\right\rangle=\omega^{m}\left|{m}\right\rangle$ and $X\left|{m}\right\rangle=\left|{m-1\,(\text{mod}\quad d)}\right\rangle$. The experimental realization of single-qudit operations can be achieved through microwave driving resonant with specific qudit state levels. For $d=3$ (i.e., a qutrit), any single-qutrit gate can be decomposed into a sequence of operations acting on two-level subspaces [104, 145], and implemented with high fidelity by accounting for the AC Stark effect arising from anharmonicity [104, 148]. It is convenient to introduce a graphical representation [Fig. 1(a)] for the generalized Pauli operators and their commutation relations. With this representation, the stabilizer $O_{p}$ is described by a loop operator around the plaquette $p$, which satisfies $O_{p}^{d}=1$ and has eigenvalues $1$, $\omega$,…, $\omega^{d-1}$. Stabilizers on different plaquettes commute with each other; therefore, the Hamiltonian Eq. (1) is exactly solvable. The ground state $\left|{\Psi_{0}}\right\rangle$ satisfies:
@@
The model defined by Eq. (1) provides a Hamiltonian realization of the topological gauge theory with gauge group $\mathbb{Z}_{d}$. The stabilizers on light (dark) plaquettes correspond to local Gauss’s law (flatness condition) constraints. The ground state, defined by Eq. (2), is the state where all local constraints are satisfied, while excited states arise from violations of these constraints. For $d=2$, the stabilizer has only two eigenvalues $1$ and $-1$, indicating whether a local constraint is satisfied or violated. The elementary excitations—violations of Gauss’s law and flatness condition—are termed charge ($e$) and flux ($m$) excitations, respectively, following its physical meaning in gauge theory. In the $\mathbb{Z}_{2}$ gauge theory, these excitations are self-dual, satisfying $e^{2}=1$ and $m^{2}=1$. For $d\geq 3$, however, charge and flux excitations are no longer self-dual and can appear in multiple powers. Specifically, the elementary excitations of Eq. (1) include charge excitations $e,e^{2},...,e^{d-1}$ on light plaquettes and flux excitations $m,m^{2},...,m^{d-1}$ on dark plaquettes (see Appendix. A for an algebraic characterization of these excitations). More general excitations involve composites $e^{k}m^{d-k}$, known as dyons. When a charge excitation $e$ encircles a flux excitation $m$ counterclockwise, or vice versa, the system acquires a phase factor $\overline{\omega}\equiv\omega^{d-1}$ due to the Aharonov-Bohm effect. Consequently, charge, flux, and dyon excitations are Abelian anyons, created by string operators [graphically represented in Fig. 1(a)]. The ground state of Eq. (1) exhibits the Abelian topological order characterized by the quantum double $D(\mathbb{Z}_{d})$ [39].